Odwa borrowed R500 from Luyanda four months ago and 700 nine months ago. He wants to liquidate these obligations with a single payment in 12 months time from now. If a simple interest rate of 11%per year, is charged on all the months, the amount he will oay luyanda in 12 months time is?

1
Expert's answer
2022-03-01T18:49:43-0500

"Amount (A)=(P\times R\times T)+P"


"A=500\times.11\times\frac{16}{12}+500=573.33"


"A=700\times.11\times\frac{21}{12}+700=834.75"


Total single payment=573.33+834.75=1,408.08